\begin{thebibliography}{101}

\bibitem[A1]{A1} T. Adachi, {\it The classification of $\tau$-tilting modules over Nakayama
algebras}, J. Algebra, 452(2016), 227-262.

\bibitem[A2]{A2} T. Adachi, {\it Characterizing $\tau$-rigid-finite algebras with radical
square zero}, Proc. Amer. Math. Soc., 144 (2016), 4673-4685.

\bibitem[Ao]{Ao} T. Aoki, {\it Classifying torsion classes for algebras with radical square zero via sign decomposition},
arXiv: 1803.03795.

%\bibitem[AAC]{AAC} T. Adachi, T. Aihara and A. Chan, {\it Tilting Brauer graph algebras
%I: Classification of two-term tilting complexes,} arXiv:1504.04827.

\bibitem[AIR]{AIR} T. Adachi, O. Iyama and I. Reiten, {\it $\tau$-tilting
theory}, Compos. Math., { 150}(3) 2014, 415-452.

%\bibitem [AiI]{AiI} T. Aihara and O. Iyama, {\it Silting mutation in
%triangulated categories}, J. Lond. Math. Soc., {85}(3) (2012),
%633-668.

\bibitem [AnHK]{AnHK} L. Angeleri H\"{u}gel, D, Happel and H. Krause,
{\it Handbook of tilting modules}, London Math. Soc. Lecture Notes
Series, {332}.

%\bibitem [AnMV]{AnMV} L. Angeleri H\"{u}gel, F. Marks, J. Vit\'oria, {\it Silting modules}, Int. Math. Res. Not. IMRN 2016, no. 4, 1251--1284.

\bibitem[AsSS]{AsSS} I. Assem, D. Simson and A. Skowro\'nski,
{\it Elements of the Representation Theory of Associative Algebras.
Vol. 1. Techniques of Reperesentation Theory,} London Math. Soc.
Student Texts, {65}, Cambridge Univ. Press, Cambridge, 2006.

%\bibitem[AuB]{AuB} M. Auslander, M. Bridger, {\it Stable module theory}, Memoirs
%Amer. Math. Soc., No. {94}, American Mathematical Society,
%Providence, R.I. 1969, 146.

%\bibitem [AuPR] {AuPR} M. Auslander, M. I. Platzeck and I. Reiten,
%{\it Coxeter functions without diagrams,} Trans. Amer. Math. Soc., {
%250} (1979), 1-12.

%\bibitem[AuR1] {AuR1} M. Auslander and I. Reiten, {\it Cohen-Macaulay and Gorenstein algebras,} In: Representation theory of finite
%groups and finite-dimensional algebras, Bielefeld, 1991, edited by
%G. O. Michler and C. M. Ringel, Progr. Math., Vol. 95, Birkhauser,
%Basel, 1991, 221-245.
%\bibitem[AuR1]{AuR1} M. Auslander and I. Reiten, {\it Applications of contravariantly finite
%subcategories,} Adv. Math., {86}(1991), 111-152.

%\bibitem[AuR2]{AuR2} M. Auslander and I. Reiten, {\it Homologically finite subcategories,} in:
%Proceedings of the Tsukuba conference, 1990, Representations of
%Algebras, London Mathematical Society Lecture Notes Series Vol 168
%(Cambridge University Press) 1-42.

%\bibitem[AuR3]{AuR3} M. Auslander and I. Reiten, {\it k-Gorenstein algebra and syzygy
%modules,} J. Pure Appl. Algebra, {92}(1994), 1-27.

%\bibitem[AuR4]{AuR4} M. Auslander and I. Reiten, {\it Syzygy modules for Neotherian
%Rings,} J. Algebra, {183}(1996), 167-185.


\bibitem[AuRS]{AuRS} M. Auslander, I. Reiten and S. O. Smal$\phi$, {\it Representation Theory of Artin
Algebras}, Cambridge Studies in Advanced Mathematics, 36. Cambridge
University Press, Cambridge, 1997.


%\bibitem[AuS1]{AuS1} M. Auslander and S. O. Smal$\phi$, {\it Preprojective over Atrin
%algebras}, J. Algebra, 66(1980), 61-122.


%\bibitem[AuS2]{AuS2} M. Auslander and S. O. Smal$\phi$, {\it Almost split sequences in
%subcategories,} J. Algebra, {69}(1981), 426-454, Addendum; J.
%Algebra, {71}(1981), 592-597.

\bibitem[B]{B} K. Bongartz, {\it Tilted Algebras}, Proc. ICRA III (Puebla 1980), Lecture Notes in Math. No.{ 903}, Springer-Verlag
1981, 26-38.

\bibitem[BGP]{BGP} I. N. Bernstein, I. M. Gelfand and V. A. Ponomarev, {\it Coxeter functors and Gabriel's theorem}, Russ. Math.
Surv. {28} (1973), 17-32.

%\bibitem[BjB]{BjB} A. Bjorner and F. Brenti, {\it Combinatorics of Coxeter groups: Graduate
%Texts in Mathematics}, 231. Springer, New York, 2005.

\bibitem[BrB]{BrB} S. Brenner and M. C. R. Butler, Generalization of the
Bernstein-Gelfand-Ponomarev reflection functors, Lecture Notes in
Math. 839, Springer-Verlag (1980), 103-169.

\bibitem[BHRR]{BHRR} T. Br\"{u}stle, L. Hille, C. M. Ringel and G. R\"{o}hrle.
 {\it The $\Delta$-Filtered Modules Without Self-Extensions for
the Auslander Algebra of $k[T]/ \langle T^{n} \rangle$,}
 Algebr. Represent. Theory, {2}(1999), 295-312.

%\bibitem[BIRS]{BIRS} A. B. Buan, O. Iyama, I. Reiten and J. Scott, {\it Cluster structures for
%2-Calabi-Yau categories and unipotent groups}. Compos. Math., { 145}
%(4)(2009), 1035-1079.

%\bibitem[BMRRT]{BMRRT} A. B. Buan, R. Marsh, M. Reineke, I. Reiten and G.
%Todorov, {\it Tilting theory and cluster combinatorics}, Adv. Math.,
%204 (2006), no. 2, 572-618.

\bibitem [C]{C} X. Chen, {\it Algebras with radical square zero are either self-injective or CM-free},
Proc. Amer. Math. Soc., 140(1)(2012), 93-98.

\bibitem[Ci]{Ci} C. Cibils, {\it Hochschild cohomology alebra of radical square zero algebras}, Eighth International
Conference on Representations of Algebras, Idun Reiten, Sverre O. Smalo, Oyvind Solberg, Aug 1996,
Geirangen, Norway, 93-101.

%\bibitem[C]{C} X. Chen, {\it Extensions of covariantly finite
%subcategories}, Arch. Math., 93(2009), 29-35.

\bibitem[CX]{CX} H. Chen and C. Xi, {\it Dominant dimensions, derived equivalences and tilting
modules}, Isr. J. Math., 215(2016), 349-395.

%\bibitem[Cl]{Cl} J. Clark, {\it Auslander-Gorenstein rings for beginners}, Trends Math.,
%International Symposium on Ring Theory, Kyongju, 1999, Birkhauser,
%Boston (2001), 95-115.

\bibitem[CHU]{CHU}
F. Coelho, D. Happel, L. Unger, {\it Complements to partial tilting modules},
J. Algebra 170 (1994), no. 1, 184--205.

\bibitem[CrS]{CrS} W. Crawley-Bovey and J. Sauter, {\it On quiver Grassmannians and orbit closures for representation-finite algebras}, Math. Z. 285 (1-2)(2017), 367--395.

\bibitem[DIJ]{DIJ} L. Demonet, O. Iyama and G. Jasso, {\it $\tau$-tilting finite algebras, bricks and $g$-vectors},  Int. Math. Res. Not., 3(2019), 852-892.

%\bibitem[DIRRT]{DIRRT} L. Demonet, O. Iyama, N. Reading, I. Reiten and H. Thomas, {\it Lattice theory of torsion classes}, arXiv:1711.01785.

%\bibitem[EJR]{EJR} F. Eisele, G. Janssens and T. Raedschelders, {\it A reduction theorem for $\tau$-rigid
%modules}, arXiv:1603.04293.

%\bibitem[FGR]{FGR} R.M. Fossum, P.A. Griffith and I.
%Reiten. \newblock{\em {T}rivial {E}xtensions of {A}belian
%{C}ategories.} Lect. Notes in Math. {456}, (Springer-Verlag,
%Berlin), 1975.

%\bibitem[Ha]{Ha} D. Happel, {\it Triangulated categories in the representation theory
%of finite-dimensional algebras,} London Mathematical Society Lecture
%Note Series, 119. Cambridge University Press, Cambridge, 1988.

\bibitem[HaR]{HaR} D. Happel and C. M. Ringel, {\it Tilted algebras,}
Trans. Amer. Math. Soc., {274} (1982), 399-443.

%\bibitem[HaU1]{HaU1} D. Happel and L. Unger, {\it Modules of finite projective dimension and
%cocovers}, Math Ann. 306 (1996), 445-457.

%\bibitem[HaU2]{HaU2} D. Happel and L. Unger, {\it On a partial order of tilting modules,}
%Algebras Represent. Theory, {8}(2005), 147-156.

%\bibitem[HaU3]{HaU3} D. Happel and L. Unger, {\it On the quiver of tilting
%modules}, J. Algebra, 284(2005), 857-868.

%\bibitem[Hu]{Hu}Z. Huang, {\it Syzygy modules for quasi $k$-Gorenstein
%algebras}, J. Algebra, 299(2006),21-32.

%\bibitem[HuI]{HuI} Z. Huang and O. Iyama, {\it Auslander type conditions and cotorsion
%pairs,} J. Algebra, 318(2007), 93-110.

%\bibitem[HuZ]{HuZ} Z. Huang and Y. Zhang, {\it G-stable support $\tau$-tilting
%modules,} Front. Math. China, 11(4)(2016), 1057-1077.

%\bibitem[I]{I} O. Iyama, {\it Cluster tilting for Higher Auslander
%algebras,} Adv. Math, 226(1) (2011), 1-61.

\bibitem[IJY]{IJY} O. Iyama, P. Jorgensen and D. Yang, {\it Intermediate
co-t-structures, two-term silting objects, $\tau$-tilting modules
and torsion classes,} Algebra Number Theory, {8}(10)(2014),
2413-2431.


%\bibitem[IY]{IY} O. Iyama and Y. Yoshino, {\it Mutations in triangulated categories and rigid Cohen-Macaulay
%modules,} Invent. Math., {172}(2008), 117-168.

\bibitem[IZ1]{IZ1} O. Iyama and X. Zhang, {\it Classifying $\tau$-tilting modules over the Auslander algebra of
$K[x]/(x^n)$}, J. Math Soc. Japan, 72(3)(2020), 731-764.

\bibitem[IZ2]{IZ2} O. Iyama and X. Zhang, {\it Tilting modules over
Auslander-Gorenstein algebra}, Pacific J. Math., 298(2)(2019), 399-416.


\bibitem[IwS]{IwS} Y. Iwanaga and H. Sato, {\it On Auslander's $n$-Gorenstein
rings,} J. Pure Appl. Algebra, 106(1996), 61-76.

%\bibitem[KR]{KR} B. Keller and I. Reiten, {\it Cluster-tilted algebras are Gorenstein
%and stably Calabi-Yau,} Adv. Math., {211}(1)(2007), 123-151.

\bibitem[J]{J} G. Jasso, {\it Reduction of $\tau$-Tilting Modules and Torsion Pairs,} Int. Math. Res. Not. IMRN,
{16}(2015), 7190-7237.

\bibitem[Je]{Je} J. Jeuenich, {\it Tilting modules over the Auslander alegebra of $K[x]/(x^n)$}, arXiV: 1803.10707.

\bibitem[K]{K} N. Kajita, {\it The number of tilting modules over
hereditary algebras and tilting modules over Auslander algebras}
Master Thesis in Graduate School of Mathematics in Nagoya
University, 2008 (in Japanese).

%\bibitem[Ma]{Ma} R. Marczinzik, {\it Auslander-Gorenstein
%algebras, standardly stratified algebras and dominant dimensions},
%arXiv:1610.02966.

%\bibitem[Mi]{M} Y. Miyashita, {\it Tilting modules of finite projective
%dimension,} Math. Z, 193(1986), 113-146.

\bibitem[Mi]{Mi} Y. Mizuno, {\it Classifying $\tau$-tilting modules over preprojective algebras
of Dynkin type,} Math. Zeit., {277}(3) (2014), 665-690.


%\bibitem[RS]{RS} C. Riedtmann, A. Schofield, {\it On a simplicial complex associated
%with tilting modules,} Comment. Math. Helv., {66}(1) (1991), 70-78.


\bibitem[NRTZ]{NRTZ} V. C. Nguyen, I. Reiten, G. Todorov, S. Zhu, {\it Dominant dimension and tilting
modules}, Math. Zeit., 292(3-4)(2019), 947-973.

\bibitem[PS]{PS} M. Pressland and J. Sauter, {\it Special tilting modules for algebras with positive dominant
dimension}, arXiv:1705.03367.


\bibitem[Ts]{Ts} Y. Tsujioka, {\it Tilting modules over the Auslander algebra of $K[x]/(x^n)$,}
Master Thesis in Graduate School of Mathematics in Nagoya
University, 2008.

\bibitem[W]{W} Z. Wang, {\it Tate-Hochschild Cohomology of Radical Square Zero Algebras}, Algebr. Represent. Th., 12(2018),1-24.

\bibitem[We]{We} J. Wei, {\it $\tau$-tilting modules and $*$-modules},
J. Algebra, 414(2014), 1-5.

\bibitem[Z1]{Z1} X. Zhang,  {\it $\tau$-rigid modules over Auslander algebras}, Taiwanese J. Math.,
21(4)(2017), 727-738.

\bibitem[Z2]{Z2} X. Zhang, {\it $\tau$-rigid modules for algebras with radical square zero}, to appear in Algebr. Colloq, arXiv:1211.5622.

%\bibitem[Zh]{Zh} Y. Zhang,  {\it On mutation of $\tau$-tilting
%modules}, Comm. Algebra, 45(6)(2017), 2726-2729.

\bibitem[Zi1]{Zi1} S. Zito, {\it $\tau$-rigid modules from tilted to cluster-tilted algebras}, Comm. Algebra, 47(9)(2019), 3716-3734.

\bibitem[Zi2]{Zi2} S. Zito, {\it $\tau$-tilting finite tilted and cluster-tilted algebras}, arXiv:1902.05866v1.



\end{thebibliography}

\documentclass[a4paper,reqno]{amsart}

\usepackage{amssymb}
\usepackage{amstext}
\usepackage{amsmath}
\usepackage{amscd}
\usepackage{amsthm}
\usepackage{amsfonts}
\usepackage{enumerate}
\usepackage{graphicx}
\usepackage{latexsym}
\usepackage{mathrsfs}
\usepackage{mathtools}
\usepackage[all]{xy}
\xyoption{all}

\usepackage{pstricks}
\usepackage{lscape}
\usepackage{comment}

\newtheorem{theorem}{Theorem}[section]

\newtheorem{proposition}[theorem]{Proposition}
\newtheorem{definition-proposition}[theorem]{Definition-Proposition}

\theoremstyle{definition}
\newtheorem{definition}[theorem]{Definition}

\newtheorem{remark}[theorem]{Remark}
\newtheorem{example}[theorem]{Example}

\newcommand{\Ext}{\operatorname{Ext}\nolimits}

\newcommand{\Hom}{\operatorname{Hom}\nolimits}
\newcommand{\End}{\operatorname{End}\nolimits}

\renewcommand{\mod}{\mathsf{mod}\hspace{.01in}}

\newcommand{\add}{\mathsf{add}\hspace{.01in}}

\newcommand{\tilt}{\mbox{\rm tilt}\hspace{.01in}}
\newcommand{\sttilt}{\mbox{\rm s$\tau$-tilt}\hspace{.01in}}

\newcommand{\gldim}{\mathop{{\rm gl.dim}}\hspace{.01in}}
\newcommand{\RHom}{\mathbf{R}\strut\kern-.2em\operatorname{Hom}\nolimits}

\numberwithin{equation}{section}

\usepackage{paralist}

\def\id{\mathop{\rm id}\nolimits}
\def\pd{\mathop{\rm pd}\nolimits}
\def\add{\mathop{\rm add}\nolimits}

\def\gldim{\mathop{\rm gldim}\nolimits}

\begin{document}
\title{Classifying tilting modules over the Auslander algebras of Radical square zero Nakayama algebras}
\thanks{2000 Mathematics Subject Classification: 16G10,16E10}
\thanks{Keywords: Auslander algebra, tilting module, support $\tau$-tilting module }
 \thanks{ The author is supported by NSFC
(Nos. 11571164 and 11671174)}

\author{Xiaojin Zhang}
\address{X. Zhang: 1. School of Mathematics and Statistics, Jiangsu Normal University, Xuzhou, 221116, P. R. China.}
\email{xjzhang@jsnu.edu.cn, xjzhangmaths@163.com}
\maketitle
\begin{abstract} Let $\Lambda$ be a radical square zero Nakayama algebra with $n$ simple modules and let $\Gamma$ be the Auslander algebra of $\Lambda$. Then every indecomposable direct summand of a tilting $\Gamma$-module is either simple or projective. Moreover, if $\Lambda$ is self-injective, then the number of tilting $\Gamma$-modules is $2^n$; otherwise, the number of tilting $\Gamma$-modules is $2^{n-1}$.
\end{abstract}

\section{Introduction}

Tilting theory has been very essential in the representation theory of finite dimensional algebras since 1970S (see \cite{B,BGP,BrB,HaR}and so on). The most important notion of tilting theory is tilting module. So it is interesting to classify tilting modules over a given algebra. For the recent development of this topics, we refer to \cite {AsSS,AnHK,CX,CrS, NRTZ, PS, IZ1}.

Auslander algebras are also important in the representation theory of finite dimensional algebras \cite{AuRS}. Tilting modules over Auslander algebras have got a lot of attention. In 2000, Br\"{u}stle and his coauthors \cite{BHRR} studied the tilting modules over the Auslander algebra of $K[x]/(x^n)$ and showed that the number of tilting modules is $n!$. Indeed, this algebra is an Auslander algebra of a Nakayama local algebra. In 2008, Kajita \cite{K} studied the number of tilting modules over the Auslander algebra of hereditary algebra of Dynkin type in his master's thesis. In 2014, Adachi, Iyama and Reiten \cite{AIR} introduced the notions of $\tau$-tilting modules as generalizations of tilting modules. For more details of $\tau$-tilting modules, we refer to \cite{DIJ,IJY,J,Mi,We,Z1,Zi1,Zi2} and the references there. In 2016, Iyama and Zhang \cite{IZ1,Ts} revisit the tilting modules over the Auslander algebra of $K[x]/(x^n)$ and built a bijection between the tilting modules over the Auslander algebra of $K[x]/(x^n)$ and the support $\tau$-tilting modules over the corresponding preprojective algebra of $A_{n-1}$ \cite {Mi}. All algebras studied in \cite{K,BHRR,IZ1} are the Auslander algebras of some Nakayama algebras. For an arbitrary Auslander algebra $\Lambda$, finding all the tilting $\Lambda$-modules seems difficult.

On the other hand, radical square zero algebras have gained a lot of attention in the representation theory of finite-dimensional algebras. Auslander, Reiten and Smal$\phi$ \cite{AuRS} showed that these algebras are stable equivalent to hereditary algebras. Chen \cite{C} showed that these algebras are either self-injective or CM-free (all finitely generated Gorenstein projective modules are projective). The Hochschild cohomology algebra of radical square zero algebras are studied by Cibils \cite{Ci}. In 2018, Wang \cite {W} studied the Tate-Hochschild cohomology of radical square zero algebras. Moreover, $\tau$-tilting modules over these algebras have studied by Adachi \cite{A2} and the author \cite{Z2}. Aoki \cite{Ao} classified torsion classes over algebras of radical square zero. It is showed by Auslander that there is a bijection between Auslander algebras and algebras of finite representation type. It is natural to consider the Auslander algebras of radical square zero algebras of finite representation type.

In this paper, we focus on tilting modules over the Auslander algebras of radical square zero Nakayama algebras and show the following theorem.
\begin{theorem} Let $\Lambda$ be a radical square zero Nakayama algebra with $n$ simple modules and let $\Gamma$ be the Auslander algebra of $\Lambda$. If $\Lambda$ is self-injective, then the number of tilting $\Gamma$-modules is $2^n$; otherwise, the number of tilting $\Gamma$-modules is $2^{n-1}$.
\end{theorem}

We show the organization of this paper as follows:
In Section 2, we recall some preliminaries and show the main result.

Throughout this paper, all algebras are finite-dimensional algebras and all modules are finitely generated right modules. For a tilting module, we mean the classical tilting module. We use $\tau$ to denote the Auslander-Reitein translation functor.

\bigskip
\noindent{\bf Acknowledgement} This note can be seen as an application of the published result proved by Iyama and the author. The author would like to thank Iyama for useful suggestions and allowing him to write this note alone.

\section{Main result and its proof}

In this section, we firstly recall some preliminaries on tilting modules, $\tau$-tilting modules and Auslander algebras. Then we prove the main result.

For an algebra $\Lambda$, denote by $\gldim\Lambda$ the global dimension of $\Lambda$ and by $I^i(\Lambda)$ the $(i+1)$-th term of minimal injective presentation of $\Lambda$. We have the following definition of Auslander algebras (resp. Auslander $1$-Gorenstein algebras).

\begin{definition}\label{2.2} An algebra $\Lambda$ is called an Auslander algebra if $\gldim\Lambda\leq2$ and $I^0(\Lambda), I^1(\Lambda)$ are projective. An algebra $\Lambda$ is called an Auslander $1$-Gorenstein if $I^0(\Lambda)$ is projective (\cite{IwS}).
\end{definition}

Recall that an algebra $R$ is of finite representation type if $\mod R=\add M$. Now we can give the definition of Auslander algebras.
Let $R$ be an algebra of finite representation type and $M$ be the additive generator of $\mod R$. We call $\Lambda=\End_R M$ the Auslander algebra of $R$.

For any $M\in\mod\Lambda$, denote by $\pd_\Lambda M$ (resp. $\id_\Lambda M$ ) the projective dimension (resp. injective dimension) of $M$. Denote by $|M|$ the number of non-isomorphic indecomposable direct summand of $M$. Now we recall the definition of tilting modules.

\begin{definition}\label{2.3} Let $\Lambda$ be an algebra and $T\in \mod\Lambda$. $T$ is called a tilting module if the following are satisfied:
\begin{itemize}
\item[\rm(1)] $\pd_\Lambda T\leq 1$

\item[\rm(2)] $\Ext_{\Lambda}^i(T,T)=0$ for $i\geq 1$.

\item[\rm(3)] $|T|=|\Lambda|$.
\end{itemize}
\end{definition}

In the following we recall the definition of $\tau$-tilting modules introduced in \cite{AIR}.

\begin{definition}\label{2.4} Let $\Lambda$  be an algebra and $M\in\mod\Lambda$. $M$ is called $\tau$-rigid if $\Hom_{\Lambda}(M,\tau M)=0$. $M$ is $\tau$-tilting if $M$ is $\tau$-rigid and $|M|=|\Lambda|$. Moreover $M$ is call a support $\tau$-tilting module if it is a $\tau$-tilting module over $\Lambda/(e)$, where $e$ is an idempotent of $\Lambda$.
\end{definition}

The following results in \cite{IZ2} shows the connection between the tilting modules over Auslander 1-Gorenstein algebras $\Lambda$ and support $\tau$-tilting modules over the factor algebra of $\Lambda$.

\begin{theorem}\label{2.5} Let $\Lambda$ be an Auslander $1$-Gorenstein algebra and $e
$ be an idempotent such that $e\Lambda$ is the additive generator of projective-injective modules. Then there is a bijection between the set
$\tilt\Lambda$ of tilting modules over $\Lambda$ and the set $\sttilt\Lambda/(e)$ of support $\tau$-tilting modules over $\Lambda/(e)$.
\end{theorem}





 Let $\Gamma$ be the Auslander algebra of a radical square zero Nakayama algebra $\Lambda$ with $n$ simple modules. Now we need to show the structure of $\Gamma$.

 \begin{proposition}\label{3.1} Let $\Gamma$ be the Auslander algebra of a radical square zero Nakayama algebra $\Lambda$ with $n$ simple modules.
 \begin{itemize}
\item[\rm(1)] If $\Lambda$ is not self-injective, then $\Gamma$ is given by the quiver $Q_1:$ $$\xymatrix{1&2\ar[l]_{a_2}&3\ar[l]_{a_3}&\cdots \ar[l]&2n-2\ar[l]_{a_{2n-2}}&2n-1\ar[l]_{a_{2n-1}}}$$ with the relations: $a_{2k-1}a_{2k-2}=0$ for $1\leq k\leq n$;
\item[\rm(2)] If $\Lambda$ is self-injective, then $\Gamma$ is given by the quiver $Q_2:$
$$\xymatrix{
&2n\ar[ld]_{a_{2n}}&1\ar[l]_{a_1}&\\
2n-1\ar[d]_{a_{2n-1}}&&&2\ar[ul]_{a_2}\\
2n-2&&&3\ar[u]_{a_3}\\
&5\ar@{.}[lu]\ar[r]_{a_5}&4\ar[ur]_{a_4}& }$$
with the relations: $a_{2k}a_{2k-1}=0$ for $1\leq k\leq n$.
\end{itemize}
\end{proposition}

\begin{proof} By using straight calculation, one gets the quiver of $\Gamma$.
\end{proof}

 In the rest of this section, we assume that $\Lambda$ and $\Gamma$ are as in Proposition \ref{3.1}.

 For a module $T\in \mod \Lambda$, denote by $Gen(T)$ the subcategory of $\mod\Lambda$ generated by $T$. For two tilting modules $T_1$ and $T_2$, we call $T_1\leq T_2$ if $Gen(T_1)\subseteq Gen(T_2)$. Now we show the properties of tilting modules over $\Gamma$.

\begin{proposition}\label{3.2} Let $T$ be an arbitrary tilting module over $\Gamma$.
 \begin{itemize}
 \item [\rm(1)] Then any indecomposable direct summand $M$ of $T$ is either projective or a simple socle of an indecomposable projective-injective $\Gamma$-module.
 \item[\rm(2)] For any indecomposable projective and non-injective direct summand $P$ of $T$, the mutation over $P$ is given by the exact sequence $0\rightarrow P\rightarrow I^0({\rm soc}P)\rightarrow S\rightarrow0$ with $S$ simple.

 \item[\rm (3)] The minimal tilting module is a direct sum of projective-injective modules and simple modules of projective dimension $1$.
 \end{itemize}
\end{proposition}

\begin{proof} (1) By \cite[Chapter V, Theorem 3.2]{AsSS}, $\Gamma$ is a Nakayama algebra with Loewy length at most $3$.
By \cite[Lemma 3.2] {Z1}, every tilting module $T$ can be a submodule of a projective module. And hence any indecomposable direct summand $M$ should be a submodule of a projective module. Since $\Gamma$ is an Auslander algebra, $M$ can be embedded into a projective-injective module by \cite[Proposition 10]{IwS}. Since $\Gamma$ is a Nakayama algebra Loewy length at most $3$, then it is not difficult to show that $M$ is either projective or the simple socle of an indecomposable projective-injective modules.

(2) Since $T\setminus P$ is faithful, then the mutation over $P$ always exists by \cite[Proposition 1.3]{CHU}. And hence we get an exact sequence $0\rightarrow P\rightarrow T_1\rightarrow U\rightarrow 0$ with $U$ indecomposable. If $P$ is simple projective, then one gets the assertion since $\Gamma$ is a Nakayama algebra. Otherwise, the length of $P$ should be $2$ by Proposition \ref{3.1}. Then the top of $P$ is of projective dimension $2$ which is not an indecomposable direct summand of $T$. Then by (1) we get $\Hom_\Gamma(P, T')=0$ for any $T'\not=I^0({\rm soc}P)$. And therefore the minimal left $\add{(T\setminus P)}$ of $P$ should be $0\rightarrow P\rightarrow I^0({\rm soc}P)\rightarrow S\rightarrow0$.

(3) By \cite[Theorem 3.4]{IZ2}, one gets that the minimal tilting module according to the partial order defined above should be $I^0(\Gamma)\oplus \Omega^{-1}\Gamma$. By Proposition 3.1, $\Gamma$ is an Auslander algebra with $\Omega^{-1}\Gamma$ semi-simple. So the assertion holds.
\end{proof}

To show the main result, we need the following observation.

\begin{proposition}\label{3.3} Let $A$ be a semi-simple algebra with $n$ simple modules. Then the number of support $\tau$-tilting $A$-modules is $2^n$.
\end{proposition}
\begin{proof} We only need to show that any direct sum of non-isomorphic simple modules is a support $\tau$-tilting $A$-module. Let $m$ be an arbitrary integer with $1\leq m\leq n$ and $S_i$ be a simple $A$-module for $1\leq i\leq n$. Then $\oplus_{i=1}^m S_i$ is also projective since $A$ is semi-simple. And hence $\oplus_{i=1}^m S_i$ is $\tau$-rigid. Let $e$ be the idempotent $1-\Sigma_{i=1}^m e_i$ with $e_i$ the idempotent according to the simple module $S_i$. Then $\oplus_{i=1}^m S_i$ is a $\tau$-tilting (in fact tilting module) over $A/(e)$. Notice that the module $0$ is support $\tau$-tilting, then the number of support $\tau$-tilting modules over $A$ should be $\Sigma_{m=0}^{n} C_{n}^{m}=2^n$. The assertion holds.
\end{proof}

For a vertex $i$ in a quiver $Q$, we denote by $P(i)$, $I(i)$ and $S(i)$ the indecomposable projective, injective and simple module according to the $i$, respectively. Now we are ready to show the following main result.

\begin{theorem}\label{3.4}Let $\Lambda$ be a radical square zero Nakayama algebra with $n$ simple modules and let $\Gamma$ be the Auslander algebra of $\Lambda$. If $\Lambda$ is self-injective, then the number of tilting $\Gamma$-modules is $2^n$; otherwise, the number of tilting $\Gamma$-modules is $2^{n-1}$.
\end{theorem}

\begin{proof}
(1) The case of $\Lambda$ is self-injective.
By Proposition \ref{3.1}, one gets the indecomposable projective-injective $\Gamma$-modules are as follows:
$P(2n)=I(2n-1), P(2n-1)=I(2n-3), \cdots, P(3)=I(1), P(1)=I(2n-1)$. Take the idempotent $e=e_{2n}+e_{2n-1}+\cdots+e_1$. Then the factor algebra $\Gamma/(e)$ is a semi-simple algebra with $n$ simple modules. By Theorem \ref{2.5} and Proposition \ref{3.3}, we get the number of tilting modules over $\Gamma$ is $2^n$.

(2) The case of $\Lambda$ is non-self-injective.

By Proposition \ref{3.1}, one gets the indecomposable projective-injective modules are as follows:
$P(2n-1)=I(2n-2),P(2n-2)=I(2n-4),\cdots, P(4)=I(2), P(2)=I(1)$. Take the idempotent $e=e_{2n-1}+e_{2n-2}+\cdots+e_2$. Then the factor algebra $\Gamma/(e)$ is semi-simple with $n-1$ simple modules. Then by Theorem \ref{2.5} and Proposition \ref{3.3}, one gets that the number of tilting $\Gamma$-modules is $2^{n-1}$.
\end{proof}

Now we give an example to show our main results:

\begin{example} Let $\Lambda$  and $\Gamma$ be as in Proposition \ref{3.1} and let $n=3$.
\begin{itemize}
\item[\rm (1)] If $\Lambda$ is non-self-injective, then the tilting $\Gamma$-modules are as follows:

$T_1=\Gamma, T_2=P(5)\oplus P(4)\oplus S(4)\oplus P(2)\oplus P(1)$

$ T_3=P(5)\oplus P(4)\oplus P(3)\oplus P(2)\oplus S(2), T_4=P(5)\oplus P(4)\oplus S(4)\oplus P(2)\oplus S(2)$.

\item[\rm (2)] If $\Lambda$ is self-injective, then the tilting $\Gamma$-modules are as follows:

$T_1=\Gamma, T_2=S(1)\oplus P(5)\oplus P(4)\oplus P(3)\oplus P(2)\oplus P(1)$;

$T_3=P(6)\oplus P(5)\oplus S(5)\oplus P(3)\oplus P(2)\oplus P(1), T_4=P(6)\oplus P(5)\oplus P(4)\oplus P(3)\oplus S(3)\oplus P(1)$;

$T_5=S(1)\oplus P(5)\oplus S(5)\oplus P(3)\oplus P(2)\oplus P(1), T_6=P(6)\oplus P(5)\oplus S(5)\oplus P(3)\oplus S(3)\oplus P(1)$;

$T_7=S(1)\oplus P(5)\oplus P(4)\oplus P(3)\oplus S(3)\oplus P(1), T_8=S(1)\oplus P(5)\oplus S(5)\oplus P(3)\oplus S(3)\oplus P(1)$.

\end{itemize}
\end{example}

We end the paper with the following remarks.

\begin{remark}
\begin{itemize}

\item[\rm (1)] Let $\Gamma$ be the Auslander algebra of $K[x]/(x^2)$. Then it is showed in \cite{BHRR, IZ1} that the number of tilting modules over $\Gamma$ should be $2$ which is equal to $2^1$. So Theorem \ref{3.4} is also true.
\item [\rm (2)]Tilting modules of finite projective dimension over the the Auslander algebra of $K[x]/(x^n)$ have been classified by Geuenich \cite{Je}. However, the classification of tilting modules of finite projective dimension over the Auslander algebras of radical square zero Nakayama algebras is still open.
\item [\rm (3)] Adachi in \cite{A1} classified $\tau$-tilting modules over Nakayama algebras in terms of triangulations. It is also interesting to find the correspondence between tilting modules and triangulations.

\end{itemize}
\end{remark}

\end{document}